# Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction

### v1

@@ Abstract
Aspect sentiment quadruple prediction (ASQP) facilitates a detailed understanding of opinions expressed in a text by identifying the opinion term, aspect term, aspect category and sentiment polarity for each opinion. However, annotating a full set of training examples to fine-tune models for ASQP is a resource-intensive process. In this study, we explore the capabilities of large language models (LLMs) for zero- and few-shot learning on the ASQP task across five diverse datasets. We report F1 scores slightly below those obtained with state-of-the-art fine-tuned models but exceeding previously reported zero- and few-shot performance. In the 40-shot setting on the Rest16 restaurant domain dataset, LLMs achieved an F1 score of 52.46, compared to 60.39 by the best-performing fine-tuned method MVP. Additionally, we report the performance of LLMs in target aspect sentiment detection (TASD), where the F1 scores were also close to fine-tuned models, achieving 66.03 on Rest16 in the 40-shot setting, compared to 72.76 with MVP. While human annotators remain essential for achieving optimal performance, LLMs can reduce the need for extensive manual annotation in ASQP tasks.
@@ E.1 TASD
Aspect Term - 0 63.26 53.03 78.38 69.02 61.16 79.22 63.97 54.61 77.23 65.00 57.62 74.55 64.01 54.29 77.99
10 70.00 66.15 74.35 73.68 72.05 75.52 74.41 70.16 79.21 65.74 65.68 65.83 67.89 63.54 72.91
20 71.37 66.72 76.72 75.72 74.06 77.48 75.21 68.27 83.72 67.70 64.47 71.47 69.46 63.11 77.25
30 71.15 67.40 75.35 76.32 73.72 79.15 74.91 68.26 83.01 69.64 65.53 74.31 70.35 65.30 76.25
40 71.07 68.95 73.32 75.86 76.36 75.39 74.46 68.77 81.19 70.62 66.04 75.88 70.49 67.11 74.23
50 70.59 68.54 72.77 75.63 75.35 75.95 74.93 70.01 80.59 70.19 64.62 76.82 71.62 68.22 75.38
SC 0 64.98 55.66 78.04 71.23 65.35 78.27 64.99 57.78 74.26 66.38 60.08 74.17 65.48 57.79 75.51
10 72.63 71.84 73.43 76.21 79.09 73.53 75.54 74.66 76.44 67.82 70.98 64.93 67.88 67.57 68.19
20 72.43 70.77 74.17 78.66 80.62 76.80 75.49 71.30 80.20 69.24 68.36 70.14 72.32 70.59 74.14
30 72.29 71.90 72.69 78.68 79.27 78.10 76.05 71.83 80.79 73.60 74.22 72.99 71.13 73.01 69.34
40 73.63 75.19 72.14 76.42 82.42 71.24 76.38 73.58 79.41 73.47 72.79 74.17 72.62 74.64 70.71
50 73.26 74.81 71.77 76.45 80.99 72.39 76.82 75.29 78.42 74.07 72.40 75.83 71.93 74.21 69.79
Aspect Category - 0 60.30 60.60 60.00 76.05 76.08 76.02 67.04 66.06 68.06 77.11 75.60 78.68 65.77 65.13 66.44
10 71.67 74.95 68.67 76.90 78.56 75.32 75.27 75.81 74.74 75.41 77.73 73.24 66.43 68.89 64.14
20 73.72 76.99 70.71 80.45 81.08 79.84 76.92 77.78 76.09 77.25 76.67 77.84 68.02 69.00 67.08
30 75.24 78.36 72.36 81.84 82.62 81.08 78.97 79.81 78.16 79.84 79.81 79.88 71.19 73.10 69.38
40 76.86 80.05 73.91 81.44 82.85 80.08 80.29 80.86 79.74 80.63 80.03 81.24 71.86 74.49 69.42
50 77.15 80.82 73.81 80.69 81.70 79.70 79.83 80.46 79.21 80.58 79.14 82.08 71.69 74.12 69.42
SC 0 60.17 61.85 58.58 75.52 77.67 73.49 67.22 68.97 65.55 76.69 76.39 77.00 66.09 68.58 63.77
10 72.43 81.12 65.42 78.27 86.11 71.74 76.67 81.11 72.70 75.35 82.42 69.40 67.53 75.46 61.10
20 73.71 81.76 67.10 81.63 87.04 76.85 78.61 83.44 74.31 77.68 82.00 73.80 69.68 76.09 64.27
30 74.93 84.22 67.48 82.38 88.06 77.39 80.55 85.06 76.50 80.93 86.04 76.40 73.15 83.19 65.28
40 78.15 87.14 70.84 82.76 90.03 76.58 82.04 86.93 77.66 81.39 85.81 77.40 72.98 82.22 65.61
50 78.68 88.67 70.71 82.52 89.47 76.58 81.85 86.89 77.37 81.85 86.09 78.00 73.87 82.64 66.78
Sentiment Polarity - 0 89.49 90.01 88.98 90.34 89.58 91.11 93.29 93.97 92.62 91.24 90.60 91.88 90.47 90.48 90.47
10 90.07 91.91 88.30 89.81 91.48 88.19 94.05 95.24 92.89 89.69 91.81 87.68 88.26 92.06 84.76
20 90.95 92.32 89.63 92.19 92.96 91.43 94.19 95.57 92.85 91.83 91.92 91.74 91.59 92.21 90.99
30 91.23 92.27 90.22 92.53 93.26 91.81 94.63 95.90 93.40 92.43 92.05 92.81 91.79 93.15 90.47
40 91.13 92.45 89.85 92.62 93.47 91.78 94.43 95.21 93.68 93.06 92.72 93.40 91.54 92.53 90.57
50 90.51 92.17 88.92 92.82 93.53 92.13 94.93 95.79 94.09 92.86 91.66 94.08 92.94 94.36 91.56
SC 0 88.54 91.15 86.07 88.78 90.31 87.30 91.36 95.35 87.69 90.26 91.04 89.49 88.37 91.86 85.14
10 87.07 94.39 80.80 86.94 94.76 80.32 91.45 96.58 86.84 87.61 93.85 82.15 85.49 94.83 77.83
20 87.53 93.81 82.04 89.56 94.53 85.08 91.46 96.40 87.01 89.32 94.29 84.84 89.28 94.71 84.43
30 88.25 94.84 82.51 89.06 94.00 84.60 93.14 97.21 89.40 90.70 94.68 87.04 87.56 94.78 81.37
40 88.03 94.34 82.51 89.54 95.50 84.29 92.60 96.83 88.72 90.56 94.67 86.80 89.53 94.97 84.67
50 87.40 94.11 81.58 89.90 95.70 84.76 93.13 97.39 89.23 90.59 94.43 87.04 89.81 96.23 84.20
@@
Aspect Term - 0 51.80 47.39 57.23 54.53 53.23 55.98 49.94 45.89 54.77 57.86 56.47 59.72 52.79 53.45 52.22
10 56.98 53.66 61.00 61.41 58.23 65.07 65.75 64.94 66.61 49.48 56.99 43.79 52.03 57.22 47.78
20 61.04 57.29 65.39 63.21 61.08 65.59 67.36 61.36 74.69 53.83 55.18 52.56 54.10 50.46 58.40
30 63.05 59.72 66.79 66.51 65.71 67.35 67.95 61.06 76.63 56.92 55.01 59.00 56.37 54.36 58.58
40 63.97 62.27 65.79 67.00 65.76 68.30 67.63 63.55 72.40 60.03 57.01 63.46 56.64 55.27 58.12
50 63.68 61.94 65.57 67.42 64.96 70.16 68.13 64.28 72.51 57.83 52.95 63.79 58.59 57.15 60.14
SC 0 51.70 51.56 51.85 53.44 55.56 51.47 52.05 51.25 52.87 58.81 61.72 56.16 55.25 60.88 50.57
10 56.07 76.43 44.28 61.19 82.32 48.69 64.20 80.78 53.27 48.51 79.89 34.83 49.93 71.19 38.44
20 67.25 83.15 56.46 67.58 83.61 56.70 69.49 74.72 64.95 56.04 75.50 44.55 56.51 79.92 43.71
30 66.16 81.40 55.72 69.41 86.76 57.84 71.26 75.44 67.52 60.50 79.54 48.82 58.52 77.15 47.14
40 67.67 81.35 57.93 69.47 83.83 59.31 70.65 78.31 64.36 66.21 78.76 57.11 57.14 76.05 45.77
50 67.94 79.65 59.23 70.72 82.71 61.76 71.00 76.91 65.94 64.62 74.01 57.35 59.55 77.09 48.51
Aspect Category - 0 41.03 46.58 36.70 60.47 67.13 55.05 46.14 50.82 42.25 73.39 79.74 68.20 53.53 65.45 45.31
10 55.53 54.03 57.16 55.70 51.49 60.70 64.72 64.20 65.26 59.92 64.48 55.96 54.34 58.22 50.95
20 60.05 59.24 60.90 62.61 56.89 69.64 70.26 69.55 71.01 64.95 64.21 65.72 50.65 49.57 51.79
30 60.89 59.77 62.06 67.93 63.33 73.24 68.58 65.25 72.29 69.49 67.66 71.44 56.29 55.72 56.89
40 64.12 62.83 65.47 66.81 63.73 70.20 73.01 70.19 76.09 70.68 68.22 73.36 58.69 59.15 58.26
50 63.83 63.94 63.74 67.86 65.03 70.96 73.14 70.69 75.77 70.80 67.31 74.72 59.14 59.68 58.63
SC 0 37.11 45.99 31.10 59.72 69.98 52.09 43.55 51.08 37.96 72.81 84.43 64.00 52.66 70.39 42.07
10 53.69 80.05 40.39 54.24 77.50 41.72 66.73 85.78 54.60 56.56 89.22 41.40 58.53 82.87 45.24
20 60.61 82.96 47.74 64.56 78.05 55.05 72.25 85.92 62.34 63.38 85.96 50.20 55.01 82.61 41.24
30 58.90 78.96 46.97 68.70 85.09 57.60 71.43 82.85 62.77 66.00 88.55 52.60 60.17 83.88 46.91
40 63.08 79.84 52.13 66.29 81.82 55.72 75.35 87.31 66.28 70.11 87.69 58.40 59.30 81.58 46.58
50 65.45 84.49 53.42 67.31 82.87 56.66 75.54 84.48 68.32 69.95 84.66 59.60 62.05 85.84 48.58
Sentiment Polarity - 0 85.30 90.82 80.43 86.10 90.06 82.48 90.97 94.98 87.28 89.32 91.27 87.48 89.48 93.67 85.66
10 88.31 89.87 86.81 87.18 87.35 87.02 88.58 89.41 87.76 86.28 89.06 83.67 85.94 91.99 80.66
20 88.52 89.94 87.15 88.17 86.72 89.68 89.16 89.34 88.99 84.35 83.31 85.43 90.69 91.85 89.58
30 87.70 88.12 87.28 88.05 86.55 89.62 91.31 91.11 91.52 84.61 83.27 86.01 90.67 92.29 89.10
40 88.82 89.54 88.11 88.77 87.51 90.06 89.92 89.21 90.67 87.27 85.80 88.80 90.97 92.12 89.86
50 87.81 89.33 86.35 89.86 88.74 91.02 90.50 89.71 91.32 87.87 85.66 90.22 91.46 92.78 90.19
SC 0 81.69 91.88 73.53 84.54 91.50 78.57 87.74 95.20 81.37 87.92 92.68 83.62 85.60 95.10 77.83
10 68.00 96.58 52.48 67.77 96.48 52.22 76.92 96.15 64.10 67.19 95.50 51.83 76.79 97.81 63.21
20 72.76 94.55 59.13 77.29 95.13 65.08 83.41 96.41 73.50 74.85 94.76 61.86 74.45 97.70 60.14
30 74.10 96.29 60.22 78.12 95.63 66.03 84.47 96.90 74.87 75.52 96.93 61.86 78.09 96.53 65.57
40 78.31 96.38 65.94 78.39 95.44 66.51 84.50 96.70 75.04 79.20 95.83 67.48 79.05 96.92 66.75
50 76.98 95.00 64.71 79.33 96.58 67.30 87.03 96.66 79.15 83.08 97.06 72.62 79.72 97.94 67.22

### v2

@@ Abstract
Aspect sentiment quad prediction (ASQP) facilitates a detailed understanding of opinions expressed in a text by identifying the opinion term, aspect term, aspect category and sentiment polarity for each opinion. However, annotating a full set of training examples to fine-tune models for ASQP is a resource-intensive process. In this study, we explore the capabilities of large language models (LLMs) for zero- and few-shot learning on the ASQP task across five diverse datasets. We report F1 scores slightly below those obtained with state-of-the-art fine-tuned models but exceeding previously reported zero- and few-shot performance. In the 40-shot setting on the Rest16 restaurant domain dataset, LLMs achieved an F1 score of 52.46, compared to 60.39 by the best-performing fine-tuned method MVP. Additionally, we report the performance of LLMs in target aspect sentiment detection (TASD), where the F1 scores were also close to fine-tuned models, achieving 66.03 on Rest16 in the 40-shot setting, compared to 72.76 with MVP. While human annotators remain essential for achieving optimal performance, LLMs can reduce the need for extensive manual annotation in ASQP tasks.
@@ E.1 TASD
Aspect Term - 0 63.26 53.03 78.38 69.02 61.16 79.22 65.00 57.62 74.55 51.30 39.17 74.30 64.01 54.29 77.99
10 70.00 66.15 74.35 73.68 72.05 75.52 65.74 65.68 65.83 62.91 53.38 76.63 67.89 63.54 72.91
20 71.37 66.72 76.72 75.72 74.06 77.48 67.70 64.47 71.47 66.58 56.91 80.27 69.46 63.11 77.25
30 71.15 67.40 75.35 76.32 73.72 79.15 69.64 65.53 74.31 67.18 57.15 81.51 70.35 65.30 76.25
40 71.07 68.95 73.32 75.86 76.36 75.39 70.62 66.04 75.88 68.70 60.57 79.38 70.49 67.11 74.23
50 70.59 68.54 72.77 75.63 75.35 75.95 70.19 64.62 76.82 69.30 60.89 80.41 71.62 68.22 75.38
SC 0 64.98 55.66 78.04 71.23 65.35 78.27 66.38 60.08 74.17 54.66 44.26 71.48 65.48 57.79 75.51
10 72.63 71.84 73.43 76.21 79.09 73.53 67.82 70.98 64.93 63.47 61.15 65.98 67.88 67.57 68.19
20 72.43 70.77 74.17 78.66 80.62 76.80 69.24 68.36 70.14 69.54 67.09 72.16 72.32 70.59 74.14
30 72.29 71.90 72.69 78.68 79.27 78.10 73.60 74.22 72.99 72.61 71.05 74.23 71.13 73.01 69.34
40 73.63 75.19 72.14 76.42 82.42 71.24 73.47 72.79 74.17 71.67 71.19 72.16 72.62 74.64 70.71
50 73.26 74.81 71.77 76.45 80.99 72.39 74.07 72.40 75.83 73.38 72.88 73.88 71.93 74.21 69.79
Aspect Category - 0 60.30 60.60 60.00 76.05 76.08 76.02 77.11 75.60 78.68 42.36 39.63 45.50 65.77 65.13 66.44
10 71.67 74.95 68.67 76.90 78.56 75.32 75.41 77.73 73.24 40.49 39.84 41.16 66.43 68.89 64.14
20 73.72 76.99 70.71 80.45 81.08 79.84 77.25 76.67 77.84 45.25 44.54 45.99 68.02 69.00 67.08
30 75.24 78.36 72.36 81.84 82.62 81.08 79.84 79.81 79.88 46.07 45.47 46.69 71.19 73.10 69.38
40 76.86 80.05 73.91 81.44 82.85 80.08 80.63 80.03 81.24 48.00 47.99 48.02 71.86 74.49 69.42
50 77.15 80.82 73.81 80.69 81.70 79.70 80.58 79.14 82.08 48.42 48.25 48.60 71.69 74.12 69.42
SC 0 60.17 61.85 58.58 75.52 77.67 73.49 76.69 76.39 77.00 42.47 43.01 41.94 66.09 68.58 63.77
10 72.43 81.12 65.42 78.27 86.11 71.74 75.35 82.42 69.40 43.92 51.24 38.43 67.53 75.46 61.10
20 73.71 81.76 67.10 81.63 87.04 76.85 77.68 82.00 73.80 47.79 54.81 42.36 69.68 76.09 64.27
30 74.93 84.22 67.48 82.38 88.06 77.39 80.93 86.04 76.40 49.58 58.22 43.18 73.15 83.19 65.28
40 78.15 87.14 70.84 82.76 90.03 76.58 81.39 85.81 77.40 49.47 57.18 43.60 72.98 82.22 65.61
50 78.68 88.67 70.71 82.52 89.47 76.58 81.85 86.09 78.00 52.01 60.77 45.45 73.87 82.64 66.78
Sentiment Polarity - 0 89.49 90.01 88.98 90.34 89.58 91.11 91.24 90.60 91.88 89.00 87.29 90.78 90.47 90.48 90.47
10 90.07 91.91 88.30 89.81 91.48 88.19 89.69 91.81 87.68 86.45 86.07 86.83 88.26 92.06 84.76
20 90.95 92.32 89.63 92.19 92.96 91.43 91.83 91.92 91.74 88.68 88.53 88.83 91.59 92.21 90.99
30 91.23 92.27 90.22 92.53 93.26 91.81 92.43 92.05 92.81 89.07 88.54 89.61 91.79 93.15 90.47
40 91.13 92.45 89.85 92.62 93.47 91.78 93.06 92.72 93.40 89.27 89.03 89.51 91.54 92.53 90.57
50 90.51 92.17 88.92 92.82 93.53 92.13 92.86 91.66 94.08 89.86 89.35 90.39 92.94 94.36 91.56
SC 0 88.54 91.15 86.07 88.78 90.31 87.30 90.26 91.04 89.49 84.84 88.80 81.22 88.37 91.86 85.14
10 87.07 94.39 80.80 86.94 94.76 80.32 87.61 93.85 82.15 79.84 90.43 71.46 85.49 94.83 77.83
20 87.53 93.81 82.04 89.56 94.53 85.08 89.32 94.29 84.84 81.13 90.66 73.41 89.28 94.71 84.43
30 88.25 94.84 82.51 89.06 94.00 84.60 90.70 94.68 87.04 81.82 93.99 72.44 87.56 94.78 81.37
40 88.03 94.34 82.51 89.54 95.50 84.29 90.56 94.67 86.80 83.45 93.09 75.61 89.53 94.97 84.67
50 87.40 94.11 81.58 89.90 95.70 84.76 90.59 94.43 87.04 82.61 93.25 74.15 89.81 96.23 84.20
@@
Aspect Term - 0 51.80 47.39 57.23 54.53 53.23 55.98 57.86 56.47 59.72 30.09 26.14 35.46 52.79 53.45 52.22
10 56.98 53.66 61.00 61.41 58.23 65.07 49.48 56.99 43.79 43.69 39.85 48.73 52.03 57.22 47.78
20 61.04 57.29 65.39 63.21 61.08 65.59 53.83 55.18 52.56 52.66 48.03 58.35 54.10 50.46 58.40
30 63.05 59.72 66.79 66.51 65.71 67.35 56.92 55.01 59.00 53.45 48.25 60.14 56.37 54.36 58.58
40 63.97 62.27 65.79 67.00 65.76 68.30 60.03 57.01 63.46 59.10 54.13 65.22 56.64 55.27 58.12
50 63.68 61.94 65.57 67.42 64.96 70.16 57.83 52.95 63.79 56.36 51.85 61.79 58.59 57.15 60.14
SC 0 51.70 51.56 51.85 53.44 55.56 51.47 58.81 61.72 56.16 32.47 30.77 34.36 55.25 60.88 50.57
10 56.07 76.43 44.28 61.19 82.32 48.69 48.51 79.89 34.83 31.23 77.03 19.59 49.93 71.19 38.44
20 67.25 83.15 56.46 67.58 83.61 56.70 56.04 75.50 44.55 40.41 79.00 27.15 56.51 79.92 43.71
30 66.16 81.40 55.72 69.41 86.76 57.84 60.50 79.54 48.82 44.84 83.96 30.58 58.52 77.15 47.14
40 67.67 81.35 57.93 69.47 83.83 59.31 66.21 78.76 57.11 55.56 85.11 41.24 57.14 76.05 45.77
50 67.94 79.65 59.23 70.72 82.71 61.76 64.62 74.01 57.35 53.05 83.70 38.83 59.55 77.09 48.51
Aspect Category - 0 41.03 46.58 36.70 60.47 67.13 55.05 73.39 79.74 68.20 43.10 48.73 38.64 53.53 65.45 45.31
10 55.53 54.03 57.16 55.70 51.49 60.70 59.92 64.48 55.96 25.76 24.75 26.86 54.34 58.22 50.95
20 60.05 59.24 60.90 62.61 56.89 69.64 64.95 64.21 65.72 28.54 26.81 30.54 50.65 49.57 51.79
30 60.89 59.77 62.06 67.93 63.33 73.24 69.49 67.66 71.44 28.05 25.70 30.91 56.29 55.72 56.89
40 64.12 62.83 65.47 66.81 63.73 70.20 70.68 68.22 73.36 31.91 29.29 35.08 58.69 59.15 58.26
50 63.83 63.94 63.74 67.86 65.03 70.96 70.80 67.31 74.72 31.01 28.57 33.93 59.14 59.68 58.63
SC 0 37.11 45.99 31.10 59.72 69.98 52.09 72.81 84.43 64.00 42.08 52.47 35.12 52.66 70.39 42.07
10 53.69 80.05 40.39 54.24 77.50 41.72 56.56 89.22 41.40 15.46 37.90 9.71 58.53 82.87 45.24
20 60.61 82.96 47.74 64.56 78.05 55.05 63.38 85.96 50.20 20.92 44.90 13.64 55.01 82.61 41.24
30 58.90 78.96 46.97 68.70 85.09 57.60 66.00 88.55 52.60 25.04 47.95 16.94 60.17 83.88 46.91
40 63.08 79.84 52.13 66.29 81.82 55.72 70.11 87.69 58.40 28.28 46.89 20.25 59.30 81.58 46.58
50 65.45 84.49 53.42 67.31 82.87 56.66 69.95 84.66 59.60 25.32 41.71 18.18 62.05 85.84 48.58
Sentiment Polarity - 0 85.30 90.82 80.43 86.10 90.06 82.48 89.32 91.27 87.48 83.65 88.73 79.12 89.48 93.67 85.66
10 88.31 89.87 86.81 87.18 87.35 87.02 86.28 89.06 83.67 86.26 84.46 88.15 85.94 91.99 80.66
20 88.52 89.94 87.15 88.17 86.72 89.68 84.35 83.31 85.43 88.81 88.18 89.46 90.69 91.85 89.58
30 87.70 88.12 87.28 88.05 86.55 89.62 84.61 83.27 86.01 87.02 85.10 89.02 90.67 92.29 89.10
40 88.82 89.54 88.11 88.77 87.51 90.06 87.27 85.80 88.80 87.85 87.00 88.73 90.97 92.12 89.86
50 87.81 89.33 86.35 89.86 88.74 91.02 87.87 85.66 90.22 88.35 87.60 89.12 91.46 92.78 90.19
SC 0 81.69 91.88 73.53 84.54 91.50 78.57 87.92 92.68 83.62 80.00 91.25 71.22 85.60 95.10 77.83
10 68.00 96.58 52.48 67.77 96.48 52.22 67.19 95.50 51.83 41.97 93.28 27.07 76.79 97.81 63.21
20 72.76 94.55 59.13 77.29 95.13 65.08 74.85 94.76 61.86 47.46 92.25 31.95 74.45 97.70 60.14
30 74.10 96.29 60.22 78.12 95.63 66.03 75.52 96.93 61.86 53.40 93.87 37.32 78.09 96.53 65.57
40 78.31 96.38 65.94 78.39 95.44 66.51 79.20 95.83 67.48 59.63 95.19 43.41 79.05 96.92 66.75
50 76.98 95.00 64.71 79.33 96.58 67.30 83.08 97.06 72.62 61.49 95.38 45.37 79.72 97.94 67.22